If R1000 and the average fixed cost is R16 when 25 units of output are produced, then the average variable cost at that level of output is...

1. R16
2. R24
3. R40
4. Impossible to determine.
1
Expert's answer
2018-03-05T08:33:07-0500
So, average fixed cost is R16 for 1 units of output, that`s why average fixed cost for all units is R16*25 = R400, then average variable cost for all units is R1000 – R400 = R600, so average variable cost for 1 units is R600/25 = R24. So Answer is R24(2.)